# Test-Time Alignment of Discrete Diffusion Models with Sequential Monte Carlo

### v1

@@ 2.1 Discrete Diffusion Models
Discrete diffusion models (Austin et al., 2021) have a forward process which interpolates between the original data distribution and a fixed prior $\boldsymbol{\pi}\in\Delta^{V}$:
@@ 3.1 Twisted SMC with Locally Optimal Proposal
Figure 1: From left to right: samples from the pre-trained distribution $p_{\mathrm{pre}}$; samples from the target distribution $p_{\mathrm{tar}}$, samples generated using tempered SMC with the locally optimal proposal (Eq. 11), the reverse process as proposal (Eq. 10), and the first order Taylor approximation of the locally optimal proposal (Sec. 3.2); samples generated using approximate guidance (Eq. 24). KL weight $\alpha=1$ for both distributions. Top row: reward $r(X,Y)=-\hat{X}^{2}/100-\hat{Y}^{2}$, bottom row: reward $r(X,Y)=-\hat{X}^{2}-(\hat{Y}-1)^{2}/10$, where $\hat{X}=12(X/63-1/2)$ and $\hat{Y}=12(Y/63-1/2)$. Note that the rewards are actually computed as a differentiable function of one hot inputs corresponding to $X$ and $Y$, more details in Sec. E.2.1.
@@
$\mathcal{F}_{s}(\mathbf{z}^{1:L}_{s}|\mathbf{z}^{1:L}_{t})=\prod_{l=1}^{L}m_{s}(\mathbf{z}^{l}_{s}|\mathbf{z}^{1:L}_{t}),$
@@ 4 Experiments
We evaluate the proposed method on both a synthetic dataset and an image modelling task. Detailed experimental settings and additional results are presented in Appendix E.
@@ 4.1 Mixture of Gaussians
Results. In Fig. 1, we provide a comparison of the resulting samples when using twisted SMC with different proposals, and when using approximate guidance only. We report the mean reward achieved, EMD (earth mover’s distance) w.r.t. the target distribution, and sample diversity (the number of unique samples). SMC with the locally optimal proposal gives the best results, achieving the required reward value while having a low EMD and high sample diversity. SMC with the first order approximation of the locally optimal proposal also achieves comparable results. When using the reverse process as the proposal, there is a significant dip in sample diversity, which is caused due the higher variance in weights, which in turn leads to more resampling steps. Lastly, the approximate guidance performs the worst as the gradient term overpowers the diffusion reverse process, leading to samples which disregard the pre-trained distribution. A further analysis on the effect of using different proposals is presented in Sec. E.2.2.
@@ 4.2 Binarised MNIST
The above stated property of masked diffusion models also makes it somewhat difficult to sample using the twisted SMC method. We needed to carefully choose the tempering schedule to get good results. Since, unmasked pixels can not be modified further, we need to ensure that a sufficient amount of guidance is injected into the proposal early on before the particles evolve into an incorrect digit. At the same time, we cannot increase the temperature too fast since the guidance can be inaccurate in earlier time steps. This means we need to find a balance. Our choice of tempering schedule for this experiment is visualised in Fig. 6. This problem is resolved if we use either ReMDM (Wang et al., 2025) or UDLM (Schiff et al., 2024) in place of masked diffusion, as they allow pixels to be modified and guided throughout the entire sampling process. We provide the details of our experiments with ReMDM and UDLM on the binarised MNIST dataset in Sec. E.3.
@@ 5 Conclusions and Limitations
In this work, we present a method for test-time reward alignment of discrete diffusion models using twisted SMC. We introduce a first-order Taylor approximation of the locally optimal proposal, and provide methods to counter the discontinuities when computing the gradients of the estimated reward for masked diffusion models. The effectiveness of our method is validated through empirical results on both synthetic and image datasets. We also discuss limitations and potential future work in Appendix F.
@@ A.2 Uniform Noise Diffusion Models
Uniform Noise Diffusion Models (UDLM) (Schiff et al., 2024) is a discrete diffusion model, where the fixed prior $\boldsymbol{\pi}=\boldsymbol{u}:={\bm{1}}/V$, where $V$ is the vocabulary size or the number of possible categorical values. The resulting posterior is,
@@
Appendix B Sampling Algorithm

### v2

@@ 2.1 Discrete Diffusion Models
Discrete diffusion models (Austin et al., 2021) define a forward process that interpolates data with prior $\boldsymbol{\pi}\in\Delta^{V}$
@@ 3.1 Twisted SMC with Locally Optimal Proposal
Figure 1: From left to right: samples from the pre-trained distribution $p_{\mathrm{pre}}$; samples from the target distribution $p_{\mathrm{tar}}$, samples generated using tempered SMC with the locally optimal proposal (Eq. 11), the reverse process as proposal (Eq. 10), and the first order Taylor approximation of the locally optimal proposal (Sec. 3.2); samples generated using approximate guidance (Eq. 39). KL weight $\alpha=1$ for both distributions. Top row: reward $r(X,Y)=-\hat{X}^{2}/100-\hat{Y}^{2}$, bottom row: reward $r(X,Y)=-\hat{X}^{2}-(\hat{Y}-1)^{2}/10$, where $\hat{X}=12(X/63-1/2)$ and $\hat{Y}=12(Y/63-1/2)$. Note that the rewards are actually computed as a differentiable function of one hot inputs corresponding to $X$ and $Y$, more details in Sec. F.2.1.
@@
$\mathcal{F}_{s}(\mathbf{z}^{1:L}_{s}|\mathbf{z}^{1:L}_{t})=\prod_{l=1}^{L}\mathcal{F}_{s}(\mathbf{z}^{l}_{s}|\mathbf{z}^{1:L}_{t}),$
@@ 4 Experiments
We evaluate the proposed method on both a synthetic dataset and an image modelling task. Detailed experimental settings and additional results are presented in Appendix F.
@@ 4.1 Mixture of Gaussians
Results. In Fig. 1, we provide a comparison of the resulting samples when using twisted SMC with different proposals, and when using approximate guidance only. We report the mean reward achieved, EMD (earth mover’s distance) w.r.t. the target distribution, and sample diversity (the number of unique samples). SMC with the locally optimal proposal gives the best results, achieving the required reward value while having a low EMD and high sample diversity. SMC with the first order approximation of the locally optimal proposal also achieves comparable results. When using the reverse process as the proposal, there is a significant dip in sample diversity, which is caused due the higher variance in weights, which in turn leads to more resampling steps. Lastly, the approximate guidance performs the worst as the gradient term overpowers the diffusion reverse process, leading to samples which disregard the pre-trained distribution. A further analysis on the effect of using different proposals is presented in Sec. F.2.2.
@@ 4.2 Binarised MNIST
The above stated property of masked diffusion models also makes it somewhat difficult to sample using the twisted SMC method. We needed to carefully choose the tempering schedule to get good results. Since, unmasked pixels can not be modified further, we need to ensure that a sufficient amount of guidance is injected into the proposal early on before the particles evolve into an incorrect digit. At the same time, we cannot increase the temperature too fast since the guidance can be inaccurate in earlier time steps. This means we need to find a balance. Our choice of tempering schedule for this experiment is visualised in Fig. 6. This problem is resolved if we use either ReMDM (Wang et al., 2025) or UDLM (Schiff et al., 2024) in place of masked diffusion, as they allow pixels to be modified and guided throughout the entire sampling process. We provide the details of our experiments with ReMDM and UDLM on the binarised MNIST dataset in Sec. F.3.
@@ 5 Conclusions and Limitations
In this work, we present a method for test-time reward alignment of discrete diffusion models using twisted SMC. We introduce a first-order Taylor approximation of the locally optimal proposal, and provide methods to counter the discontinuities when computing the gradients of the estimated reward for masked diffusion models. The effectiveness of our method is validated through empirical results on both synthetic and image datasets. We also discuss limitations and potential future work in Appendix G.
@@ A.2 Uniform Noise Diffusion Models
Uniform Noise Diffusion Models (UDLM) (Schiff et al., 2024) is a discrete diffusion model where the fixed prior $\boldsymbol{\pi}=\boldsymbol{u}:={\bm{1}}/V$, where $V$ is the vocabulary size or the number of possible categorical values. The resulting posterior is,
@@
Appendix B Extending Discrete-time SMC to Continuous-time SMC
In this section, we establish the connection between the discrete-time SMC formulation in Eq. 4 and its continuous-time counterparts, as studied in (Ou et al., 2025; Holderrieth et al., 2025; Lee et al., 2025). We begin by recapitulating continuous-time Markov chains (CTMC) (Norris, 1998) in Sec. B.1 and then discuss the connection in Sec. B.2.
B.1 Background of CTMC
A CTMC (Norris, 1998) is characterised by a time-dependent transition rate matrix $R_{t}$, defined as
$R_{t}(\mathbf{x},\mathbf{y})=\lim_{\Delta t\rightarrow 0}\frac{p_{t+\Delta t|t}(\mathbf{y}|\mathbf{x})-\delta_{\mathbf{y}=\mathbf{x}}}{\Delta t}.$ (22)
By definition, the transition probability is given by
$p_{t+\Delta t|t}(\mathbf{y}|\mathbf{x})=\delta_{\mathbf{y}=\mathbf{x}}+R_{t}(\mathbf{x},\mathbf{y})\Delta t+\mathcal{O}(\Delta t).$ (23)
To ensure the transition probability is valid, the rate matrix $R_{t}$ should satisfy
$R_{t}(\mathbf{x},\mathbf{y})\geq 0,\forall\mathbf{y}\neq\mathbf{x},\quad R_{t}(\mathbf{x},\mathbf{x})=-\sum_{\mathbf{y}\neq\mathbf{x}}R_{t}(\mathbf{x},\mathbf{y}).$ (24)
The CTMC transition probabilities satisfy the Kolmogorov forward and backward equations (Oksendal, 2013). For $t>s$
Kolmogorov forward equation: $\quad\partial_{t}q_{t|s}(\mathbf{x}|\tilde{\mathbf{x}})=\sum_{\mathbf{y}}q_{t|s}(\mathbf{y}|\tilde{\mathbf{x}})R_{t}(\mathbf{y},\mathbf{x})$
Kolmogorov backward equation: $\quad\partial_{s}q_{t|s}(\mathbf{x}|\tilde{\mathbf{x}})=-\sum_{\mathbf{y}}R_{t}(\tilde{\mathbf{x}},\mathbf{y})q_{t|s}(\mathbf{x}|\mathbf{y})$
The forward equation also induces a PDE for the marginals of CTMC
$\partial_{t}q_{t}(\mathbf{x})=\sum_{\mathbf{y}}q_{t}(\mathbf{y})R_{t}(\mathbf{y},\mathbf{x}).$ (25)
Given a test function $h$ of interest, define $u_{t}(\mathbf{x})={\mathbb{E}}_{q_{1|t}(\mathbf{z}|\mathbf{x})}[h(\mathbf{z})]$, the backward equation gives
$\partial_{t}u_{t}(\mathbf{x})$ $=\sum_{\mathbf{z}}h(\mathbf{z})\partial_{t}q_{1|t}(\mathbf{z}|\mathbf{x})$
$=\sum_{\mathbf{z}}h(\mathbf{z})-\sum_{y}R_{t}(\mathbf{x},\mathbf{y})q_{1|t}(\mathbf{z},\mathbf{y})$
$=-\sum_{y}R_{t}(\mathbf{x},\mathbf{y})\sum_{\mathbf{z}}q_{1|t}(\mathbf{z},\mathbf{y})h(\mathbf{z})$
$=-\sum_{y}R_{t}(\mathbf{x},\mathbf{y})u_{t}(\mathbf{y}).$
B.2 CTMC Formulation of SMC
Recall Eq. 4, where the importance weight is given by
$w_{s}(\mathbf{x}_{s})=w_{t}(\mathbf{x}_{t})\frac{\pi_{s}(\mathbf{x}_{s})}{\pi_{t}(\mathbf{x}_{t})}\frac{\mathcal{B}_{s}(\mathbf{x}_{t}|\mathbf{x}_{s})}{\mathcal{F}_{s}(\mathbf{x}_{s}|\mathbf{x}_{t})}.$ (26)
We now extend this formulation to the continuous-time setting. Let $R_{t}$ and $\hat{R}_{t}$ denote the rate matrices corresponding to the forward proposal $\mathcal{F}_{t}$ and the backward transition $\mathcal{B}_{t}$, respectively. Consider a discretization with $N$ denoising steps, indexed by time points $s=t_{N}<\dots<t_{i}\dots<t_{0}=1$, where each interval satisfies $t_{i-1}-t_{i}=\frac{1}{N}$. The discrete-time importance weight at step time $s$ is then computed as
$\log w_{s}=\log\frac{\pi_{s}(\mathbf{x}_{s})}{\pi_{1}(\mathbf{x}_{1})}+\sum_{i=1}^{N}\log\frac{\mathcal{B}_{t_{i}}(\mathbf{x}_{t_{i-1}}|\mathbf{x}_{t_{i}})}{\mathcal{F}_{t_{i}}(\mathbf{x}_{t_{i}}|\mathbf{x}_{t_{i-1}})}.$ (27)
The second term in the RHS can be computed as
$\log$ $\frac{\mathcal{B}_{t_{i}}(\mathbf{x}_{t_{i-1}}|\mathbf{x}_{t_{i}})}{\mathcal{F}_{t_{i}}(\mathbf{x}_{t_{i}}|\mathbf{x}_{t_{i-1}})}=\log\left(\delta_{\mathbf{x}_{t_{i\!-\!1}}\!=\!\mathbf{x}_{t_{i}}}\!\!+\hat{R}_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i-1}})\frac{1}{N}\right)\!-\!\log\left(\delta_{\mathbf{x}_{t_{i}}\!=\!\mathbf{x}_{t_{i-1}}}\!\!+R_{t_{i}}(\mathbf{x}_{t_{i\!-\!1}},\mathbf{x}_{t_{i}})\frac{1}{N}\right)$
$\!\!\!\!\!\!=\!\!\!\!\sum_{i,t_{i}=t_{i\!-\!1}}\!\!\!\!\log\!\left(\!1\!+\!\hat{R}_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i}})\frac{1}{N}\!\right)\!\!-\!\log\!\left(\!1\!+\!R_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i}})\frac{1}{N}\!\right)\!\!+\!\!\!\!\!\!\sum_{i,t_{i}\neq t_{i\!-\!1}}\!\!\!\!\!\hat{R}_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i\!-\!1}})\!-\!R_{t_{i}}(\mathbf{x}_{t_{i\!-\!1}},\mathbf{x}_{t_{i}})$
$\!\!\!\!\!\!=\!\!\!\!\sum_{i,t_{i}=t_{i\!-\!1}}\!\!\!\!\hat{R}_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i-1}})\frac{1}{N}\!-\!R_{t_{i}}(\mathbf{x}_{t_{i\!-\!1}},\mathbf{x}_{t_{i}})\frac{1}{N}+\mathcal{O}(\frac{1}{N})\!+\!\!\!\!\sum_{i,t_{i}\neq t_{i\!-\!1}}\!\!\!\!\hat{R}_{t_{i}}(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i-1}})\!-\!R_{t_{i}}(\mathbf{x}_{t_{i\!-\!1}},\mathbf{x}_{t_{i}})$
Taking the limit as $N\rightarrow+\infty$, the importance weight becomes:
$\!\!\!\log w_{s}\!=\!\log\frac{\pi_{s}(\mathbf{x}_{s})}{\pi_{1}(\mathbf{x}_{1})}\!+\!\!\int_{1}^{s}\!\!R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\!-\!\hat{R}_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\mathop{}\!\mathrm{d}t+\!\!\!\!\!\!\!\!\!\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\!\!\!\!\!\!\!\!\log\hat{R}_{t}(\mathbf{x}_{t},\mathbf{x}_{t^{+}})\!-\!\log R_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t}).$ (28)
By the fundamental theorem of calculus for piecewise differentiable functions, we also have:
$\log\frac{\pi_{s}(\mathbf{x}_{s})}{\pi_{1}(\mathbf{x}_{1})}=\int_{1}^{s}-\partial_{t}\log\pi_{t}(\mathbf{x}_{t})+\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\log\pi_{t}(\mathbf{x}_{t})-\log\pi_{t}(x_{t^{+}}).$ (29)
If the backward process is chosen such that the rate matrix satisfies the detailed balance condition: $\hat{R}_{t}(\mathbf{x},\mathbf{y})p_{t}(\mathbf{x})=R_{t}(\mathbf{y},\mathbf{x})p_{t}(\mathbf{y})$, then the importance weight simplifies accordingly
$\log w_{s}$ $=\int_{1}^{s}-\partial_{t}\log\pi_{t}(\mathbf{x}_{t})+R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\!-\!\hat{R}_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\mathop{}\!\mathrm{d}t$
$=\int_{1}^{s}-\partial_{t}\log\pi_{t}(\mathbf{x}_{t})+\sum_{\mathbf{y}}R_{t}(\mathbf{x}_{t},\mathbf{y})\frac{\pi_{t}(\mathbf{y})}{\pi_{t}(\mathbf{x}_{t})}.$
This coincides with the importance weights used in the SMC proposed in (Ou et al., 2025; Holderrieth et al., 2025). In these works, the intermediate target distribution is defined as a geometric interpolation between the base and target distributions: $\pi_{t}=p_{\mathrm{base}}^{t}p_{\mathrm{target}}^{1-t}$. The proposal rate matrix $R_{t}$ is then trained to satisfy the Kolmogorov forward equation. Alternatively, one may use an arbitrary proposal, such as a pretrained rate matrix, but this typically leads to suboptimal performance.
In Lee et al. (2025), the intermediate target is defined as
$\pi_{t}(\mathbf{x}_{t})=p_{t}(\mathbf{x}_{t})p_{t}(\zeta|\mathbf{x}_{t})^{\alpha},\quad p_{t}(\zeta|\mathbf{x}_{t})={\mathbb{E}}_{p_{t}(\mathbf{x}_{0}|\mathbf{x}_{t})}[p(\zeta|\mathbf{x}_{0})],$ (30)
where $p(\zeta|\mathbf{x}_{0})$ denotes the tilting reward function. To align with the notation used in Lee et al. (2025), we denote the proposal rate matrix by $Q_{t}$, the forward rate matrix generating $p_{t}$ by $R_{t}$, and its corresponding backward rate matrix by $\hat{R}_{t}$. With this notation, the importance weight is given by:
[…]
To connect with the derivation in Lee et al. (2025), we use the following identities:
$\hat{R}_{t}(\mathbf{x}_{t},\mathbf{x}_{t^{+}})=R_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})\frac{p_{t}(\mathbf{x}_{t^{+}})}{p_{t}(\mathbf{x}_{t})}.$ (32)
$\hat{R}_{t}(\mathbf{x}_{t},\mathbf{x}_{t})=-\sum_{y\neq\mathbf{x}_{t}}\hat{R}_{t}(\mathbf{x}_{t},\mathbf{y})=-\sum_{\mathbf{y}\neq\mathbf{x}_{t}}R(\mathbf{y},\mathbf{x}_{t})\frac{p_{t}(\mathbf{y})}{p_{t}(\mathbf{x})}.$ (33)
$\partial_{t}p_{t}(\mathbf{x}_{t})=\sum_{\mathbf{y}}R_{t}(\mathbf{y},\mathbf{x}_{t})p_{t}(\mathbf{y}).$ (34)
$\partial_{t}\log p_{t}(\mathbf{x}_{t})=\frac{1}{p_{t}(\mathbf{x}_{t})}\sum_{\mathbf{y}}R_{t}(\mathbf{y},\mathbf{x}_{t})p_{t}(\mathbf{y})=\sum_{\mathbf{y}\neq\mathbf{x}}R_{t}(\mathbf{y},\mathbf{x}_{t})\frac{p_{t}(\mathbf{y})}{p_{t}(\mathbf{x}_{t})}+R_{t}(\mathbf{x}_{t},\mathbf{x}_{t}).$ (35)
$\partial_{t}\log p_{t}(\zeta|\mathbf{x}_{t})$ $=\frac{\partial_{t}p_{t}(\zeta|\mathbf{x}_{t})}{p_{t}(\zeta|\mathbf{x}_{t})}=-\frac{1}{p_{t}(\zeta|\mathbf{x}_{t})}\sum_{\mathbf{y}}R_{t}(\mathbf{x}_{t},\mathbf{y})p_{t}(\zeta|\mathbf{y})$ (36)
$=-\sum_{\mathbf{y}\neq\mathbf{x}_{t}}R_{t}(\mathbf{x}_{t},\mathbf{y})\frac{p_{t}(\zeta|\mathbf{y})}{p_{t}(\zeta|\mathbf{x}_{t})}-R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})$
$\triangleq R_{t}^{\alpha=1}(\mathbf{x}_{t},\mathbf{x}_{t})-R_{t}(\mathbf{x}_{t},\mathbf{x}_{t}),$
where we denote $R_{t}^{\alpha=1}(\mathbf{x}_{t},\mathbf{x}_{t})=R_{t}(\mathbf{x}_{t},\mathbf{y})\frac{p_{t}(\zeta|\mathbf{y})}{p_{t}(\zeta|\mathbf{x}_{t})}$, which induces the identity:
$\log p_{t}(\zeta|\mathbf{y})-\log p_{t}(\zeta|\mathbf{x}_{t})=\log R_{t}^{\alpha=1}(\mathbf{x}_{t},\mathbf{x}_{t})-\log R_{t}(\mathbf{x}_{t},\mathbf{y}).$
Substituting into ①, we obtain:
circled-1 $=\!\int_{1}^{s}\!\!\!-\!\left(\!\alpha\partial_{t}\log p_{t}(\zeta|\mathbf{x}_{t})\!+\!\!\!\sum_{\mathbf{y}\neq\mathbf{x}}\!\!R_{t}(\mathbf{y},\mathbf{x}_{t})\frac{p_{t}(\mathbf{y})}{p_{t}(\mathbf{x}_{t})}\!+\!R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\!\right)\!\!+\!Q_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\!+\!\!\!\sum_{\mathbf{y}\neq\mathbf{x}_{t}}\!\!R(\mathbf{y},\mathbf{x}_{t})\frac{p_{t}(\mathbf{y})}{p_{t}(\mathbf{x})}\mathop{}\!\mathrm{d}t$
$=\!\int_{1}^{s}-\alpha\partial_{t}\log p_{t}(\zeta|\mathbf{x}_{t})-R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})+Q_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\mathop{}\!\mathrm{d}t$
$=\!\int_{1}^{s}\alpha\left(R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})-R_{t}^{\alpha=1}(\mathbf{x}_{t},\mathbf{x}_{t})\right)-R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})+Q_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\mathop{}\!\mathrm{d}t;$
For ②, we similarly have:
circled-2 $=\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\alpha\left(\log p_{t}(\zeta|\mathbf{x}_{t})-\log p_{t}(\zeta|\mathbf{x}_{t^{+}})\right)+\log R_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})-\log Q_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})$
$=\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\alpha\left(\log R_{t}^{\alpha=1}(\mathbf{x}_{t^{+}}|\mathbf{x}_{t})-\log R_{t}(\mathbf{x}_{t^{+}}|\mathbf{x}_{t^{+}})\right)+\log R_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})-\log Q_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t}).$
Combining both terms, the full expression for the importance weight becomes:
$\log w_{s}$ =+absentcircled-1circled-2=①+②= ① + ②
$=\int_{1}^{s}Q_{t}(\mathbf{x}_{t},\mathbf{x}_{t})-R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\mathop{}\!\mathrm{d}t+\!\!\!\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\!\!\!\!\log R_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})-\log Q_{t}(\mathbf{x}_{t^{+}},\mathbf{x}_{t})$
$\quad\!+\!\int_{1}^{s}\alpha\left(R_{t}(\mathbf{x}_{t},\mathbf{x}_{t})\!-\!R_{t}^{\alpha=1}(\mathbf{x}_{t},\mathbf{x}_{t})\right)\mathop{}\!\mathrm{d}t\!+\!\!\!\!\!\!\sum_{s\leq t,\mathbf{x}_{t^{+}}\neq\mathbf{x}_{t}}\!\!\!\!\!\!\!\alpha\left(\log R_{t}^{\alpha=1}(\mathbf{x}_{t^{+}}|\mathbf{x}_{t})\!-\!\log R_{t}(\mathbf{x}_{t^{+}}|\mathbf{x}_{t^{+}})\right),$
which recovers the expression for the importance weight in (Lee et al., 2025, Theorem 2).
Appendix C Sampling Algorithm